3.) When a battery is connected to the plates of a 3.00-µF capacitor, it stores a charge of 27.0 µC. What is the voltage of the battery?

[Refer to item #3] If the same capacitor is connected to another battery and 36.0 µC of charge is stored on the capacitor, what is the voltage of the battery?


1
Expert's answer
2021-05-13T07:19:56-0400

V= voltage of the battery, C= system capacitance , Q= charge stored by the capacitor


"Q= CV"


"V=\frac{Q}{C}= \frac{27\times 10^{-6} C}{3.0\times 10^{-6}F}"


"V= 9.0 V"


part two

"V=\frac{Q}{C}= \frac{36\times 10^{-6} C}{3.0\times 10^{-6}F}= 12V"

V= 12v